46 % of the 100 students in Emmie Xi Hem International School students said that they had been late in their first period classes. What is the lower bound of the 99% confidence interval for the proportion of students in the said school who had been late in their first period classes.


Note: Write your answers in decimal form and round it to 4 decimal places.


1
Expert's answer
2022-03-16T13:34:27-0400

Confidence intervals for a proportion are calculated using the following formula:

"\hat{p} \pm Z_{\alpha/2} \sqrt{\cfrac{\hat{p}(1-\hat{p}) } {n} }"

The lower bound: 

"\hat{p} - Z_{\alpha/2} \sqrt{\cfrac{\hat{p}(1-\hat{p}) } {n} }"


"\hat{p}=0.46-" the sample proportion,

"n=100-" the sample size,

for 99% confidence interval "Z_{\alpha/2} =2.576,"


"\hat{p} - Z_{\alpha/2} \sqrt{\cfrac{\hat{p}(1-\hat{p}) } {n} } =\\
=0.46 - 2.576\sqrt{\cfrac{0.46(1-0.46) } {100} } =\\
=0.3316."